OpenStudy (anonymous):

Help! What is the probability of rolling an even number on a die and then flipping a coin and having it land on heads?

OpenStudy (anonymous):

\[\frac{ 3 }{ 6 }\times \frac{ 1 }{2 } = \frac{ 1 }{ 4 }\]

OpenStudy (kenljw):

The total number of outcome is 6X2 by the multiplication rule the probability of even and a head is (3 + 1)/6X2=1/3

OpenStudy (whpalmer4):

1H no 2H yes 3H no 4H yes 5H no 6H yes 1T no 2T no 3T no 4T no 5T no 6T no 3 successful outcomes out of 12 possible. P = 3/12 = 1/4

OpenStudy (kenljw):

some it seems, must be (3 and 1)/12=1/4 thank you